# Isometries of 3-dimensional semi-Riemannian Lie groups

### v1

@@ 5. $\mathfrak{g}$ solvable with non-solvable Killing algebra
We consider here the case where $\mathfrak{g}$ is solvable but $\pazocal{L}$ is not solvable. Then, $\pazocal{L}$ is isomorphic to either $\mathfrak{so}(3,\mathbb{R})\oplus\mathbb{R}$ or $\mathfrak{sl}(2,\mathbb{R})\oplus\mathbb{R}$. Since $\mathfrak{L}$ has a 3-dimensional solvable subalgebra, then $\mathfrak{L}$ is necessarily $\mathfrak{sl}(2,\mathbb{R})\oplus\mathbb{R}$. We have that $\mathfrak{g}\cap\mathfrak{sl}(2,\mathbb{R})$ is isomorphic to $\mathfrak{aff}(\mathbb{R})$, since the intersection must be two dimensional subalgebra of $\mathfrak{sl}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ is not a possibility. In fact it is not difficult to see that, up to conjugacy, $\mathfrak{g}=\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$.
@@ 5.1. $U\in\mathfrak{sl}(2,\mathbb{R})$
Thus, the type of $\mathrm{iso}(U)$ is given by the value of $-\alpha^{2}-2\beta\gamma$.
@@ 5.1.1. Parabolic type: $\alpha^{2}+2\beta\gamma=0$
We immediately see from the matrix above that $U$ cannot be parabolic, which would correspond to $\mathrm{iso}(U)$ being nilpotent of nilpotency 3. Indeed, $\mathrm{iso}(U)$ as an infinitesimal Lorentz isometry preserves only one direction, which then must be the line generated by $\overline{T}$. Thus, the 2-dimensional factor $\mathsf{sl}(2,\mathbb{R})/\mathfrak{I}$ is also $\mathrm{iso}(U)$-invariant. However, a nilpotent infinitesimal Lorentz isometry does not preserve a supplementary hyperplane to its invariant direction.
@@ 5.2. $U\notin\mathfrak{sl}(2,\mathbb{R})$
and the type of isotropy is determined by the quantity $-\alpha^{2}-2\beta\gamma$.
@@ 6.1.1. Elliptic type
In this case, the isotropy action has a timelike invariant line and acts as a rotation on its spacelike orthogonal complement. Thus, the associated metric $q$ will be the sum of the Killing form plus a multiple of $e\otimes e$ where $e$ is a timelike vector (of $\kappa$). More concretely, there exists a basis $\{f_{1},f_{2},f_{3}\}$ such that $\mathrm{iso}(U)=\mathrm{ad}_{f_{1}}$ and $[f_{3},f_{1}]=-f_{2},[f_{3},f_{2}]=f_{2}$. This basis can be chosen such that $[f_{1},f_{2}]=f_{3}$. The Killing form is written as $2\kappa$ where
@@ 6.1.2. Hyperbolic type
In this case, the isotropy action has an invariant spacelike direction, and its orthogonal plane is then timelike and intersects the null cone in two lines corresponding to the two non-zero eigenvectors of $\mathrm{iso}(U)$. Thus, the associated metric $q$ will be the sum of the Killing form plus a multiple of $e\otimes e$ where $e$ is a spacelike vector (of $\kappa$). More concretely, there exists a basis $\{e_{1},e_{2},e_{3}\}$ such that $\mathrm{iso}(U)=\mathrm{ad}_{e_{1}}$ and $[e_{1},e_{2}]=e_{2}$, $[e_{1},e_{3}]=-e_{3}$. This basis can be chosen such that $[e_{2},e_{3}]=e_{1}$. (Notice that the basis here is already standard). The Killing form is written as $2\kappa$ where
@@ 6.1.3. Nilpotent type
In this case, we have a unipotent automorphism $A$ having an invariant null line, say, the line generated by $f_{1}$ and preserving its orthogonal plane spanned by $\{f_{1},f_{2}\}$. Hence, $A$ preserves forms admitting this plane as a kernel. In particular $A$ preserves the dual form $f^{3}$. If follows that $A$ preserves any quadratic form which is the sum of the Killing form and a multiple of $e^{3}\otimes e^{3}$. More concretely, there exists a basis $\{f_{1},f_{2},f_{3}\}$ such that $\mathrm{iso}(U)=\mathrm{ad}_{f_{1}}$ and $[f_{1},f_{2}]=f_{1},[f_{1},f_{3}]=f_{2}$. This basis can be chosen such that $[f_{2},f_{3}]=f_{3}$. The Killing form is written as $2\kappa$ where
@@ 6.2. Case $\mathfrak{g}=\mathfrak{so}(3)$
The isotropy representation is identified with the the usual representation of $\mathrm{SO}(3)$ on $\mathbb{R}^{3}$, in particular, all one-parameter subgroups are conjugated. Thus, the isotropy Lie algebra consists only of elliptic infinitesimal isometries. We consider the standard basis $\{e_{1},e_{2},e_{3}\}$ of $\mathfrak{so}(3)$ with commutator relations
$[e_{1},e_{2}]=e_{3},\quad[e_{2},e_{3}]=e_{1},\quad[e_{3},e_{1}]=e_{2}.$
@@ 7. Transversality vs Transitivity
Assume first that $N$ is abelian. Assume there exists a one-parameter $J$ acting freely. This means all orbits are open, and hence by connectedness, there exists exactly one orbit. Thus $X$ is identified with $J\sim\mathbb{R}$ acting by translation on itself. But any diffeomorphism of $\mathbb{R}$ commuting with translations is a translation. Now, let $n_{0}\in N$ having a non-trivial fixed point set: $X\neq Fix(n_{0})\neq\emptyset$. By analyticity, this is a discrete set. By commutativity, $N$ preserves $Fix(n_{0})$, and by connectedness of $N$, it acts trivially on it, that is, $Fix(n_{0})=Fix(N)$, contradicting our hypothesis.
@@
    $N$ acts by translation on $\mathbb{R}$. Since $N$ is normal in $L$, any element $l\in L$ acts by sending a translation to a translation. It follows that $l$ acts affinely.
@@
    Now an affine transformation of $\mathbb{R}$ is either a translation or has a fixed point where it has a non-trivial derivative, say $\lambda>1$, or $10<\lambda<$ 1. This implies, for the element $l\in L$, that its $\mathrm{Ad}$-action has $\lambda$ as eigenvalue, and so cannot be parabolic or elliptic.
@@ 7.1.1. Case where $L$ is a Heisenberg extension
7.1.2. Case $\mathfrak{N}=[\pazocal{L},\pazocal{L}]$ is $\mathbb{R}^{2}$ and $[\mathfrak{L},\mathfrak{L}]\subset\mathfrak{g}$
@@
7.1.3. Case $\mathfrak{N}=[\pazocal{L},\pazocal{L}]$ is $\mathbb{R}^{2}$ and $[\mathfrak{L},\mathfrak{L}]\not\subset\mathfrak{g}$
From Subsec. 3.2.2. We will see here that transitivity holds only in the case of nilpotent isotropy. So all the other potential cases do not have a 4-dimensional isometry group. Here $N$ is $[L,L]$, with Lie algebra $\mathfrak{N}=[\pazocal{L},\pazocal{L}]=\mathbb{R}Y\oplus\mathbb{R}T$. The ideal $\mathfrak{g}_{0}$ consists of $\mathbb{R}Y$ together with $RW$ where $W\in\mathbb{R}X\oplus\mathbb{R}Z$, satisfies that $\mathrm{ad}_{W}(\mathbb{R}Y\oplus\mathbb{R}T)\subset\mathfrak{g}$, equivalently, $[W,T]\in\mathfrak{g}$. It follows that $[W,T]\in\mathbb{R}Y$. If transitivity holds, then $\mathfrak{I}$ must be contained in $\mathfrak{Tr}=\mathbb{R}W\oplus\mathfrak{N}$.
@@ 8.1. Redutive complements
As can be seen, none of the $r_{\alpha}$ is a subalgebra since $[X,Y+U]=-U$ and $U\notin r_{\alpha}.$ Recall, as mentioned in Sec. 2.3, that it was proved in [FR06], that when the isotropy is connected all semi-Riemannian homogenous spaces are reductive. We have an example here of a transitive action of $\mathrm{Aff}^{+}(\mathbb{R})\times\mathbb{R}$ on $L/I$, seen as the product $(\mathbb{H}^{2},ds^{2})\times(\mathbb{R},-dt^{2})$, such that none of the reductive complements of $\mathfrak{I}$ is a subalgebra.
@@ 8.2. Sectional curvature and completeness
Regarding completeness, recall that all metrics on $\mathfrak{euc}(2)$ are complete, [BM08], but none have 4-dimensional Killing algebra, cf. Table 3. Concerning curvature, we encountered two metrics on $\mathrm{aff}(\mathbb{R})\oplus\mathbb{R}$ which have negative constant sectional curvature. One of them has 6-dimensional Killing algebra (see Sec. 5.2) and the other one has Killing algebra of dimension 4 (see Subsec. 3.2.1). There is a third (non-equivalent, cf. Appendix A) metric $p$ on $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$, given in that standard basis $\{X,Y,Z\}$ such that $[X,Y]=Y$, by
$p(X,X)=1,\,p(Y,Z)=1,\,p(Z,Z)=1,$
@@
The converse is, perhaps, of more interest. We now consider $\mathfrak{sol}$, with the standard basis $\{X,Y,Z\}$ such that $[X,Y]=Y,[X,Z]=-Z$, and the metric $q$ given by
$q(X,Z)=1,\,q(Y,Y)=1,$
@@ A.1. Normal forms of metric Lie algebras
$[e_{1},e_{2}]=e_{2}$
$[e_{3},e_{1}]=e_{3}$
$[e_{2},e_{3}]=e_{1}$
$\begin{array}[]{l}(e^{1})^{2}+\alpha_{1}((e^{2})^{2}+(e^{3})^{2})+2\alpha_{2}(e^{2}e^{3}),\,\,\alpha_{2}\neq\pm\alpha_{1}\\ (e^{1})^{2}+\alpha_{1}((e^{2})^{2}-(e^{3})^{2})+2\alpha_{2}(e^{2}e^{3}),\,\,\alpha_{1}\neq 0\\ (e^{1})^{2}+(e^{2})^{2}+2\alpha(e^{2}e^{3}),\,\,\alpha\neq 0\\ (e^{1})^{2}-(e^{2})^{2}+2\alpha(e^{2}e^{3}),\,\,\alpha\neq 0\\ (e^{1})^{2}+2(e^{1}e^{2})+2(e^{2}e^{3})\end{array}$
@@
$[e_{1},e_{2}]=e_{2}$
$[e_{3},e_{1}]=e_{3}$
$\begin{array}[]{l}(e^{1})^{2}+(e^{2})^{2}+2(e^{2}e^{3})+\alpha_{1}(e^{3})^{2},\,\,\alpha_{1}\neq 1\\ (e^{1})^{2}-(e^{2})^{2}+2(e^{2}e^{3})+\alpha_{2}(e^{3})^{2},\,\,\alpha_{2}\neq-1,\alpha_{2}<0\\ (e^{1})^{2}+2(e^{2}e^{3})\\ (e^{1})^{2}+(e^{2})^{2}+\varepsilon(e^{3})^{2},\,\,\varepsilon=\pm 1\\ (e^{3})^{2}+(e^{2})^{2}-(e^{1})^{2}\\ (e^{2})^{2}+(e^{3})^{2}+2(e^{1}e^{3})+2(e^{2}e^{3})\\ (e^{2})^{2}+2(e^{1}e^{3})\\ \end{array}$
@@
$[e_{1},e_{2}]=e_{2}$
$[e_{1},e_{3}]=\lambda e_{3}$
$\begin{array}[]{l}(e^{1})^{2}+(e^{2})^{2}+2(e^{2}e^{3})+\alpha_{1}(e^{3})^{2},\,\,\alpha_{1}\neq 1\\ (e^{1})^{2}-(e^{2})^{2}+2(e^{2}e^{3})+\alpha_{2}(e^{3})^{2},\,\,\alpha_{2}\neq-1\\ (e^{1})^{2}+\varepsilon(e^{3})^{2}+2(e^{2}e^{3}),\,\,\varepsilon=\pm 1\\ (e^{1})^{2}+2(e^{2}e^{3})\\ (e^{1})^{2}+(e^{2})^{2}+\varepsilon(e^{3})^{2},\,\,\varepsilon=\pm 1\\ (e^{1})^{2}-(e^{2})^{2}+(e^{3})^{2}\\ (e^{3})^{2}+(e^{2})^{2}-(e^{1})^{2}\\ (e^{2})^{2}+(e^{3})^{2}+2(e^{1}e^{3})+2(e^{2}e^{3})\\ (e^{2})^{2}+2(e^{1}e^{3})\\ (e^{3})^{2}+2(e^{1}e^{2})\\ \end{array}$

### v2

@@ 5. $\mathfrak{g}$ solvable with non-solvable Killing algebra
We consider here the case where $\mathfrak{g}$ is solvable but $\pazocal{L}$ is not solvable. Then, $\pazocal{L}$ is isomorphic to either $\mathfrak{so}(3,\mathbb{R})\oplus\mathbb{R}$ or $\mathfrak{sl}(2,\mathbb{R})\oplus\mathbb{R}$. Since $\mathfrak{L}$ has a 3-dimensional solvable subalgebra, $\mathfrak{L}$ is necessarily $\mathfrak{sl}(2,\mathbb{R})\oplus\mathbb{R}$. We have that $\mathfrak{g}\cap\mathfrak{sl}(2,\mathbb{R})$ is isomorphic to $\mathfrak{aff}(\mathbb{R})$, since the intersection must be a two dimensional subalgebra of $\mathfrak{sl}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ is not a possibility. In fact it is not difficult to see that, up to conjugacy, $\mathfrak{g}=\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$.
@@ 5.1. $U\in\mathfrak{sl}(2,\mathbb{R})$
Thus, the type of $\mathrm{iso}(U)$ is given by the value of $\alpha^{2}-2\beta\gamma$.
@@ 5.1.1. Parabolic type: $\alpha^{2}+2\beta\gamma=0$
We immediately see from the matrix above that $U$ cannot be parabolic, which would correspond to $\mathrm{iso}(U)$ being nilpotent of nilpotency 3. Indeed, $\mathrm{iso}(U)$ as an infinitesimal Lorentz isometry preserves only one direction, which then must be the line generated by $\overline{T}$. Thus, the 2-dimensional factor $\mathfrak{sl}(2,\mathbb{R})/\mathfrak{I}$ is also $\mathrm{iso}(U)$-invariant. However, a nilpotent infinitesimal Lorentz isometry does not preserve a supplementary hyperplane to its invariant direction.
@@ 5.2. $U\notin\mathfrak{sl}(2,\mathbb{R})$
and the type of isotropy is determined by the quantity $\alpha^{2}-2\beta\gamma$.
@@ 6.1.1. Elliptic type
In this case, the isotropy action has a timelike invariant line and acts as a rotation on its spacelike orthogonal complement. Thus, the associated metric $q$ will be the sum of the Killing form plus a multiple of $e\otimes e$ where $e$ is a timelike vector (of $\kappa$). More concretely, there exists a basis $\{f_{1},f_{2},f_{3}\}$ such that $\mathrm{iso}(U)=\mathrm{ad}_{f_{1}}$ and $f_{3},f_{1}]=-f_{2},[f_{3},f_{2}]=f_{2}$. This basis can be chosen such that $f_{1},f_{2}]=f_{3}$. The Killing form is written as $2\kappa$ where
@@ 6.1.2. Hyperbolic type
In this case, the isotropy action has an invariant spacelike direction, and its orthogonal plane is then timelike and intersects the null cone in two lines corresponding to the two non-zero eigenvectors of $\mathrm{iso}(U)$. Thus, the associated metric $q$ will be the sum of the Killing form plus a multiple of $e\otimes e$ where $e$ is a spacelike vector (of $\kappa$). More concretely, there exists a basis $\{e_{1},e_{2},e_{3}\}$ such that $\mathrm{iso}(U)=\mathrm{ad}_{e_{1}}$ and $e_{1},e_{2}]=e_{2}$, $e_{1},e_{3}]=-e_{3}$. This basis can be chosen such that $e_{2},e_{3}]=e_{1}$. (Notice that the basis here is already standard). The Killing form is written as $2\kappa$ where
@@ 6.1.3. Nilpotent type
In this case, we have a unipotent automorphism $A$ having an invariant null line, say, the line generated by $f_{1}$ and preserving its orthogonal plane spanned by $\{f_{1},f_{2}\}$. Hence, $A$ preserves forms admitting this plane as a kernel. In particular $A$ preserves the dual form $f^{3}$. It follows that $A$ preserves any quadratic form which is the sum of the Killing form and a multiple of $f^{3}\otimes f^{3}$. More concretely, there exists a basis $\{f_{1},f_{2},f_{3}\}$ such that $\mathrm{iso}(U)=\mathrm{ad}_{f_{1}}$ and $f_{1},f_{2}]=f_{1},[f_{1},f_{3}]=f_{2}$. This basis can be chosen such that $f_{2},f_{3}]=f_{3}$. The Killing form is written as $2\kappa$ where
@@ 6.2. Case $\mathfrak{g}=\mathfrak{so}(3)$
The isotropy representation is identified with the usual representation of $\mathrm{SO}(3)$ on $\mathbb{R}^{3}$, in particular, all one-parameter subgroups are conjugated. Thus, the isotropy Lie algebra consists only of elliptic infinitesimal isometries. We consider the standard basis $\{e_{1},e_{2},e_{3}\}$ of $\mathfrak{so}(3)$ with commutator relations
$e_{1},e_{2}]=e_{3},\quad[e_{2},e_{3}]=e_{1},\quad[e_{3},e_{1}]=e_{2}.$
@@ 7. Transversality vs Transitivity
Assume first that $N$ is abelian. Assume there exists a one-parameter subgroup $J$ acting freely. This means all orbits are open, and hence by connectedness, there exists exactly one orbit. Thus $X$ is identified with $J\sim\mathbb{R}$ acting by translation on itself. But any diffeomorphism of $\mathbb{R}$ commuting with translations is a translation. Now, let $n_{0}\in N$ having a non-trivial fixed point set: $X\neq Fix(n_{0})\neq\emptyset$. By analyticity, this is a discrete set. By commutativity, $N$ preserves $Fix(n_{0})$, and by connectedness of $N$, it acts trivially on it, that is, $Fix(n_{0})=Fix(N)$, contradicting our hypothesis.
@@
    NN acts by translation on $\mathbb{R}$. Since $N$ is normal in $L$, any element $l\in L$ acts by sending a translation to a translation. It follows that $l$ acts affinely.
@@
    Now an affine transformation of $\mathbb{R}$ is either a translation or has a fixed point where it has a non-trivial derivative, say $\lambda>1$, or $0<\lambda<1$. This implies, for the element $l\in L$, that its $\mathrm{Ad}$-action has $\lambda$ as eigenvalue, and so cannot be parabolic or elliptic.
@@ 7.1.1. Case where $L$ is a Heisenberg extension
7.1.2. Case $\mathfrak{N}=[\pazocal{L},\pazocal{L}]$ is $\mathbb{R}^{2}$ and $\mathfrak{L},\mathfrak{L}]\subset\mathfrak{g}$
@@
7.1.3. Case $\mathfrak{N}=[\pazocal{L},\pazocal{L}]$ is $\mathbb{R}^{2}$ and $\mathfrak{L},\mathfrak{L}]\not\subset\mathfrak{g}$
From Subsec. 3.2.2. We will see here that transitivity holds only in the case of nilpotent isotropy. So all the other potential cases do not have a 4-dimensional isometry group. Here $N$ is $[L,L]$, with Lie algebra $\mathfrak{N}=[\pazocal{L},\pazocal{L}]=\mathbb{R}Y\oplus\mathbb{R}T$. The ideal $\mathfrak{g}_{0}$ consists of $\mathbb{R}Y$ together with $\mathbb{R}W$ where $W\in\mathbb{R}X\oplus\mathbb{R}Z$, satisfies that $\mathrm{ad}_{W}(\mathbb{R}Y\oplus\mathbb{R}T)\subset\mathfrak{g}$, equivalently, $W,T]\in\mathfrak{g}$. It follows that $W,T]\in\mathbb{R}Y$. If transitivity holds, then $\mathfrak{I}$ must be contained in $\mathfrak{Tr}=\mathbb{R}W\oplus\mathfrak{N}$.
@@ 8.1. Reductive complements
As can be seen, none of the $r_{\alpha}$ is a subalgebra since [X,Y+U]=U[X,Y+U]=-U and $U\notin r_{\alpha}.$ Recall, as mentioned in Sec. 2.3, that it was proved in [FR06], that when the isotropy is connected all semi-Riemannian homogeneous spaces are reductive. We have an example here of a transitive action of $\mathrm{Aff}^{+}(\mathbb{R})\times\mathbb{R}$ on $L/I$, seen as the product $(\mathbb{H}^{2},ds^{2})\times(\mathbb{R},-dt^{2})$, such that none of the reductive complements of $\mathfrak{I}$ is a subalgebra.
@@ 8.2. Sectional curvature and completeness
Regarding completeness, recall that all metrics on $\mathfrak{euc}(2)$ are complete, [BM08], but none have 4-dimensional Killing algebra, cf. Table 3. Concerning curvature, we encountered two metrics on $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ which have negative constant sectional curvature. One of them has 6-dimensional Killing algebra (see Sec. 5.2) and the other one has Killing algebra of dimension 4 (see Subsec. 3.2.1). There is a third (non-equivalent, cf. Appendix A) metric $p$ on $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$, given in that standard basis $\{X,Y,Z\}$ such that $[X,Y]=Y$, by
p(X,X)=1,p(Y,Z)=1,p(Z,Z)=1,p(X,X)=1,\,p(Y,Z)=1,\,p(Z,Z)=1,
@@
The converse is, perhaps, of more interest. We now consider $\mathfrak{sol}$, with the standard basis $\{X,Y,Z\}$ such that [X,Y]=Y,[X,Z]=Z[X,Y]=Y,[X,Z]=-Z, and the metric $q$ given by
q(X,Z)=1,q(Y,Y)=1,q(X,Z)=1,\,q(Y,Y)=1,
@@ A.1. Normal forms of metric Lie algebras
$e_{1},e_{2}]=e_{2}$
$e_{3},e_{1}]=e_{3}$
$e_{2},e_{3}]=e_{1}$
$\begin{array}[]{l}(e^{1})^{2}+\tfrac{1}{2}(\alpha_{1}+\alpha_{2})((e^{2})^{2}+(e^{3})^{2})+(\alpha_{1}-\alpha_{2})(e^{2}e^{3}),\,\,\alpha_{1}\geq 1,\alpha_{2}\neq 0\vskip 3.0pt plus 1.0pt minus 1.0pt\\ (e^{1})^{2}+\tfrac{1}{2}(\alpha_{1}+\alpha_{2})((e^{2})^{2}+(e^{3})^{2})+(\alpha_{1}-\alpha_{2})(e^{2}e^{3}),\,\,\alpha_{1},\alpha_{2}<0\\ (e^{1})^{2}+\alpha_{1}((e^{2})^{2}-(e^{3})^{2})+2\alpha_{2}(e^{2}e^{3}),\,\,\alpha_{1}>0\\ (e^{1})^{2}+2\alpha(e^{2}e^{3})+\varepsilon(e^{2})^{2},\,\,\alpha\neq 0,\varepsilon=\pm 1\\ (e^{1})^{2}+2\sqrt{2}(e^{1}e^{3})+2(e^{2}e^{3})\end{array}$
@@
$e_{1},e_{2}]=e_{2}$
$e_{3},e_{1}]=e_{3}$
[…]
@@
$e_{1},e_{2}]=e_{2}$
$e_{1},e_{3}]=\lambda e_{3}$
[…]